# Interactive incremental learning of generalizable skills with local trajectory modulation

### v2

@@ I Introduction
We argue that in order to modulate skills locally, we require an underlying skill representation that 1) encodes trajectory distributions with both aleatoric and epistemic uncertainties—randomness in data and model knowledge gaps [7], and 2) allows for trajectory modulation trivially when modifications of the original demonstrations are required. To achieve this we build on the kernelized movement primitives (KMP) framework [8] (see Sec. III for a background review). Although task-parameterization of KMPs is briefly introduced in [8], the adaptation of local models has received little attention. We address the adaptation of local models by investigating when, where and how to add via-points. The result is an interactive learning framework of task-parameterized skills with local trajectory modulation (Sec. IV) that 1) permits the interactive correction of model imperfections locally, such that corrections ‘move’ with objects, 2) allows the definition of new task parameters interactively, when tasks change to require new/different objects and 3) ensures compliant interaction when extending skills beyond their initial duration by regulating stiffness based on the epistemic uncertainty, while simultaneously adding via-points. Figure 2 provides an overview of our approach. Our experimental evaluation (Sec. V) shows that our framework permits users to incrementally build on an initial model of a skill by interactively correcting errors and adding new behaviors in any task phase. We discuss the results and provide conclusions in Sec. VIVII.
@@ II-C Our approach and contribution
In Table I we compare our approach to related approaches on feature level. As shown in [5] task-parameterized approaches are more suitable for adapting to new situations than non-task-parameterized approaches like [30] or [31], including those involving multiple objects. To the best of our knowledge, we present the first task-parameterized approach allowing skill extensions with new frames and the application of local via-points, making it possible to move corrections with their corresponding objects and allowing interactive modulations, where corrections/extensions can be provided and incorporated into the policy online during execution.
@@ IV-D Uncertainty-aware skill extension in regions without demonstrations
Popular variable impedance schemes found in LfD regulate the robot stiffness by the inverse of covariance matrices [28, 23]. When the latter represent the aleatoric uncertainty, they provide an efficient way for robots to be more precise, by being stiffer, where demonstrations showed less variance, following a minimum intervention control principle [33, 28]. In the case of epistemic uncertainty, such schemes contribute to better compliance when the robot is uncertain about its actions [23]. In kernel-based methods, kernel hyperparameters are typically optimized for the training data, but their choice influences the behavior of the model in regions where data was not shown. For instance, the kernel length depends on the scale of the input domain, but it also dictates how quickly the epistemic uncertainty increases when moving away from the training data. [^2] Similar arguments can be made for other hyperparameters such as the noise variance, see [23]. Having ways to clearly distinguish between the two types of uncertainty enables users to better design variable impedance strategies. Similarly to [34], the covariance prediction of KMP can be decomposed into two terms, corresponding to aleatoric and epistemic components (see Appendix A-C for the derivation):
@@
$\underbrace{\bm{k}^{**}-\bm{k}^{*}\bm{K}^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{ep}}^{*}}+\underbrace{\bm{k}^{*}\left(\bm{K}+\bm{K}(\lambda\bm{\Sigma})^{-1}\bm{K}\right)^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{al}}^{*}}$ (4)
@@ V-A Evaluation on real robot
We evaluate our approach on a torque-controlled 7-DoF robot in an industrial scenario where an inner ring of a ball bearing needs to be transferred between two boxes (‘box 1’ and ‘box 2’) placed at different locations on a workbench. [^4] Find videos at github.com/DLR-RM/interactive-incremental-learning We provide $M=4$ demonstrations with different box positions, see Fig. 1-top. We use a time-driven representation with $\bm{s}_{h,m}=t_{h,m}/T_{m}$, where $t$ is a time step, and learn the end-effector position $\bm{\xi}_{h,m}=\bm{x}_{h,m}$. In order to easily re-scale the skill duration, we map all the inputs to the interval $[0,1]$ by dividing them by the duration of each demonstration $T_{m}$. The experiments start with $P=2$, with task parameters $\bm{b}^{(p)},\bm{A}^{(p)}$ representing the box positions and orientations, respectively. All local KMPs were initialized from GMMs with 12 components, trained on locally projected data, and $N=500$ inputs, equally spread over the input space. We chose a Matérn kernel ($\nu=5/2$) with length scale $l=0.1$ and noise variance 1.0 (see [35]). Other KMP hyperparameters were $\lambda_{1}=0.1$, $\lambda_{2}=1$, $\alpha=1$, chosen empirically. For completeness, we provide an overview of hyperparameters in Appendix A-E. In all experiments, via-points are added with $\gamma_{\Sigma}=1\times 10^{-8}$. Our algorithm successfully generalizes to novel box positions. However, when they deviate significantly from the demonstrated ones, TP-GMM [2], as well as the original TP-KMP formulation [8] fail (see Fig. 1-bottom).
@@ V-B Experiment 1: Improving generalization with interactively-defined local via-points
Since all demonstrations are given at the same height, the model has a high confidence for the $z$ coordinate in all frames, which leads to poor generalization when one of the boxes is moved to a different height (Fig. 1-bottom). If box 2 is moved to a lower height, the trade-off found by the model, from the expected vertical motions in the two frames, is not high enough to successfully move the ring out of box 1, leading to a collision. Using Algorithm 1, we have the ability to directly correct the robot and set via-points based on corrections made online. In this experiment we used an external force trigger, with threshold $\gamma_{F}=20N$. Figure 3 shows the results obtained for this experiment. With the robot running the model, a user applies a force at the robot end-effector, helping it avoid the collision and successfully move the ring out of the box. Via-points are added to the frame of box 1 and, through $\bm{b}^{(1)},\bm{A}^{(1)}$, are mapped to the global frame. With the added via-points, the robot achieves a success rate of $93\%$ across 15 box configurations (with three different heights), improved over the $60\%$ success rate of TP-GMM [2] and original TP-KMP [8] baselines.
@@ V-D Experiment 3: Incremental learning at non-demonstrated inputs
We consider a baseline where the gains are computed from the original KMP covariance (3) as $\bm{G}_{P}=(\mathrm{cov}\left[\bm{\xi}(\bm{s}^{*})\right])^{-1}$. To address numerical issues and ensure a maximum stiffness value we regularize $\mathrm{cov}\left[\bm{\xi}(\bm{s}^{*})\right]$, $\delta_{\mathrm{al}}\bm{\Sigma}^{*}_{\mathrm{al}}$, $\delta_{\mathrm{ep}}\bm{\Sigma}^{*}_{\mathrm{ep}}$ with a small scalar $1.5\times 10^{-3}$. Figure 5 shows that using our approach (5) the interaction forces after ${t=1.0}$ are negligible. Note that far from the training data, the KMP expectation (2) converges to zero, as we assume a zero mean prior (similarly to GPs [35]). With the improved compliance behaviors introduced by (5), the user is able to add via-points such that the robot moves up after the ring insertion. In this experiment we used a distance threshold with $\gamma_{\xi}=0.2m$, taking advantage of the reference trajectory going to zero to trigger the definition of via-points. The bottom-left image in Fig. 6 shows the increase in distance as via-points are added, while the plots on the right show the resulting robot motion and uncertainties. Appendix A-E5 shows the detailed stiffness profiles from Experiment 3 including a sensitivity analysis of the kernel length scales.
@@ Acknowledgments
The authors thank Thomas Eiband and Korbinian Nottensteiner for their discussions about the use cases.
@@ Appendix A Additional material
A-A Key notations
Table III summarizes the key notations used in our framework.
TABLE III: Description of key notations
@@ A-A Key notations
A-B Key acronyms
Table IV shows a glossary of important acronyms in our approach.
TABLE IV: Glossary of important acronyms
@@ A-B Key acronyms
A-C Decomposition of KMP covariance into a sum of epistemic and aleatoric terms
$\mathrm{cov}\left[\bm{\xi}(\bm{s}^{*})\right]=$
$=\bm{k}^{**}-\bm{k}^{*}\left[\bm{K}^{-1}-\bm{K}^{-1}\left(\bm{K}^{-1}+(\lambda\bm{\Sigma})^{-1}\right)^{-1}\bm{K}^{-1}\right]\bm{k}^{*\top}$
$=\bm{k}^{**}-\bm{k}^{*}\bm{K}^{-1}\bm{k}^{*\top}+\bm{k}^{*}\bm{K}^{-1}\left(\bm{K}^{-1}+(\lambda\bm{\Sigma})^{-1}\right)^{-1}\bm{K}^{-1}\bm{k}^{*\top}$
$=\underbrace{\bm{k}^{**}-\bm{k}^{*}\bm{K}^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{ep}}^{*}}+\underbrace{\bm{k}^{*}\left(\bm{K}+\bm{K}(\lambda\bm{\Sigma})^{-1}\bm{K}\right)^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{al}}^{*}},$
For illustrative purposes, we trained a KMP on a 2D dataset with input data on the interval $[0;1]$ and computed the mean and covariance using (2), (3) and (4) for the $[-0.5;1.5]$ interval, including both terms $\bm{\Sigma}_{\mathrm{ep}}^{*}$ and $\bm{\Sigma}_{\mathrm{al}}^{*}$. In Fig. 7(a) and Fig. 7(b) we plot the epistemic and aleatoric terms, respectively. Figure 7(c) shows that their sum gives the same result the original KMP formulation (3), which can be seen in Fig. 7(d).
(a) Epistemic uncertainty part of the KMP prediction $\bm{\Sigma}_{\mathrm{ep}}^{*}$ (4).
(b) Aleatoric uncertainty part of the KMP prediction $\bm{\Sigma}_{\mathrm{al}}^{*}$ (4).
(c) KMP covariance prediction as ${\mathrm{cov}[\bm{\xi}(\bm{s}^{*})]=\bm{\Sigma}_{\mathrm{ep}}^{*}+\bm{\Sigma}_{\mathrm{al}}^{*}}$ (4).
(d) Original KMP covariance prediction [8], computed from (3).
Figure 7: The proposed KMP covariance prediction (4) is computed explicitly as a sum of epistemic and aleatoric terms. Training data are shown as dotted black lines. Solid lines represent means and shaded areas are uncertainties.
A-D Demonstration recordings
We recorded four demonstrations, where the robot picks an inner ring of a ball bearing out of a box and places it inside another box, see Fig. 8 for the setup. We use different box poses for each demonstration and record the data by kinesthetically teaching a torque-controlled 7-DoF manipulator running a gravity compensation controller. Figure 9 shows the collected demonstrations, together with their local distributions from the perspective of frames 1 and 2, as well as the global frame (for the task parameters of demonstration 1).
Figure 8: Overview of the four given demonstrations, where an inner ring of a ball bearing is transferred from one box to another. The poses of the boxes differ in each demonstration. Exemplary demonstration trajectories are indicated as a red line, moving from ’Box 1’ to ’Box 2’.
Figure 9: Recorded demonstrations in local and global frames. The demonstrations are shown as a black doted line. The red lines and shaded area represent the mean and one standard deviation. The distribution in the global frame is generated from the task parameters of demonstration 1.
A-E Hyper-parameter overview
TABLE V: Hyper-parameter overview
Hyper-parameter Used in paper Validated range/options Remark
Scaling factor for KMP covariance $\alpha$ 1.0 [0.1; 1.0] -
Kernel function $k(.,.)$ matern2 matern2 / rbf -
between aleatoric and epistemic uncertainty
Table V shows the different hyperparameters used in this paper and a tested range of values to provide a guideline for choosing hyperparameters. We also explain what an increase or decrease of the individual parameter would lead to. Our proposed method has been designed to be relatively insensitive to hyperparameter choices. However, there are certain extreme scenarios where specific parameter settings can impact performance. For instance, using only one or two Gaussian functions or a kernel length of 1 might not yield optimal results. In practice, the choice of parameters involves trade-offs between smoothness and precision in the robot’s movement (owing largely to the kernel-based nature of our approach). This is evident in our findings (see Fig. 1415), which demonstrate how different parameter settings can affect the robot’s performance without necessarily impacting success or failure.
A-E1 KMP hyperparameter sensitivity analysis
We examine the impact of KMP hyperparameter choices in Section V on the generated trajectories and associated uncertainties. For a more comprehensive analysis, readers are referred to the original source [8]. In the following, when varying parameters, the remaining ones are kept to the values used in V.
Figure 10 shows the effect of $l$ and $\lambda_{1}$ in the mean trajectory. Lower values of $l$ tend to overfit the data, resulting in less smooth trajectories, an effect we can see in 10–left. A similar effect occurs when decreasing $\lambda_{1}$, which acts as a regularization term in KMP solution (see [8] for details).
Figure 10: Cartesian position $x$ from Box 1 frame (Fig.9, top-left). The different line colors correspond to different length scales $l$ (left) and $\lambda_{1}$ (right), which we compared.
In Fig. 11–left we see the effect of varying $\lambda_{2}$. Similarly to $\lambda_{1}$, decreasing $\lambda_{2}$ tends to lead to a poorer fit of the observed covariances. Figure 11–right shows the effect of the factor $\alpha$, which scales up/down the covariance to further improve the fit to the data. Note that, in Section V, we used $\alpha=1$, as done in [8].
Figure 11: Cartesian position $x$ from Box 1 frame. The different line colors correspond to different $\lambda_{2}$ (left) and re-scaling factor $\alpha$ (right), which we compared.
A-E2 Generalization success rate for different conditions of the ring-loading task
Table VI shows the individual successes and failures, as well as the success rates, for various box locations in the ring-loading task, comparing the performance of TP-GMM [2], TP-KMP [8] and our approach (TP-KMP with via-points from Experiment 1). Scenarios 5 and 15 were created by the authors, informed by their experience with scenarios where task-parameterized approaches may underperform. All the other 13 conditions were created in a user-informed manner by asking colleagues familiar with the workbench about configurations used in previous experiments and related projects. The aim of this section is to show how our contributions improve upon our baseline (TP-KMP) in various generalization tasks, therefore we did not compare to other approaches such as TP-ProMP [14] (a comparison on feature level can be found in Table I).
We treated any collisions with the environment that exceeded a pre-defined force threshold as failures. Notably, our method achieves a 93% success rate, compared to 60% with TP-GMM/TP-KMP. While our approach generally excels, there is one notable exception (Configuration #5), where a failure occurs due to a scenario that is significantly different from the demonstrations. In all demonstrations, the movement proceeds from left to right, while Configuration #5 would require the movement to be entirely reversed. By defining additional via-points, it was ultimately possible to achieve success using our approach. Nevertheless, we argue that if the task significantly deviates from what was demonstrated, a new set of demonstrations should be provided.
A-E3 Generalization of camera via-points to new camera location
Figure 12 shows the generalization of the via-points added to the camera frame in Experiment 2 to a new camera location. Notice how, for instance, the $x$ coordinate of the via-points in the global frame increases to match the new camera location, which is farther away from the robot base than that in Fig. 4.
Figure 12: Left: Via-points with respect to the camera frame are generalized to different locations in the robot workspace, since they ‘move’ with the task parameters. Right: Trajectory distribution for a camera location different from Fig. 4, showing the generalization properties of our method.
A-E4 Final model from Experiments 1–3
Figure 13 illustrates the expected robot motion after Experiments 1–3 (left) and the observed robot motion in terms of its Cartesian coordinates (right). The results demonstrate that the robot accurately follows the via-points, transitioning smoothly between them while adapting its precision to account for aleatoric uncertainty.
As the demonstrations are mapped to local frames, the robot does not need to align with these demonstrations in the global coordinate system. Instead, it is fulfilling the precision requirements of each frame model, extracted from (1), namely the new via-points. The robot remains stationary for $t>1.0$, indicating a period of high compliance. At around $t>1.2$, the robot begins to move in response to the emergence of via-points primarily along the z-axis, until it reaches the point where the camera checks the box.
Experiment 3 demonstrates that our approach allows for the compliant extension of a skill beyond its initial duration, which we argue is an important feature in LfD approaches. Particularly, since providing long demonstrations can be cumbersome, we believe that such an interactive and incremental framework is ideally suited to gradual and more intuitive skill transfer. It should be noted that just as a skill can be extended for $t>1.0$, leveraging uncertainty-aware stiffness regulation and via-point definition also allows it to be extended for $t<0.0$, such as to pick up a new object or change a tool before starting. A good practice for time-driven skills is to re-scale the input domain after via-points are added for unseen time instants, as this facilitates further skill extensions.
Overall, we believe that the capability to build on and refine existing skills, instead of re-demonstrating new skills each time even small variations occur, is a valuable feature for modern collaborative robots. The final execution of our experimental evaluation, as well as the steps leading up to it, described in Experiments 1–3, can be seen in detail in the accompanying video.
Figure 13: Final skill execution after Experiments 1–3. Left: Expected robot motion. The robot moves from box 1, passes in front of the newly introduced camera, moving to box 2 and finally moving upwards to verify the correct ring placement. Orange crosses indicate via-points. Right: Robot motion in Cartesian coordinates over time. The solid red lines shows the observed robot trajectories. Both aleatoric uncertainties are also depicted, as well as demonstrations and via-points.
A-E5 Detailed stiffness profiles for Experiment 3 with kernel length scale sensitivity analysis
A-F Analysis of the effect of $c_{1}$, $c_{2}$.
Figure 16: Stiffness values $\bm{G}_{P}$ for the $x$ dimension, before via-points are added. The different line colors correspond to different Sigmoide function parameters $c_{1}andc_{2}$.
Figure 17: Stiffness values $\bm{G}_{P}$ for the $x$ dimension, before via-points are added. The different line colors correspond to different epistemic and aleatoric scaling factors $\gamma_{ep}and\gamma_{al}$.
Figure 18 shows the stiffness profiles for changing values of $c_{1}$, $c_{2}$, using our approach (5) in the scenario with via-points after $t=1.0$. In Fig. 18–left, we observe that increasing $c_{1}$ has the effect of more rapidly reaching minimum stiffness after $t=1.0$. As discussed in IV-D, this is because it regulates how quickly the sigmoid function continuously switches between 0 and 1, which in our approach dictates the values of $w_{1}$, $w_{2}$. A more rapid switch is ensured by higher values of $c_{1}$. Figure 18–right highlights that smaller values of $c_{2}$ result in an earlier triggering of the transition. In summary, while $c_{1}$ regulates the rate of change, $c_{2}$ adjusts the set-point of the switching.
Note that, for certain choices of $c_{2}$ (in our experiments we used $c_{2}=1.5\times 10^{-3}$), lower values of $c_{1}$ generate values of $w_{1}$ which are close to, but not exactly, 1.0. This deviation slightly reduces stiffness in regions influenced by aleatoric uncertainty (e.g., $t<1.0$), as shown in Fig. 18–left. A similar effect is observed with $c_{2}$ (right), where lower $c_{2}$ values lead to increased stiffness in those regions. In our experiments we followed an approach of tuning $c_{1}$, $c_{2}$ based on the desired switching timing and adjusted $\delta_{\mathrm{al}}$ and $\delta_{\mathrm{ep}}$ to achieve the desired stiffness range for our robotic system.
Figure 18: Stiffness values $\bm{G}_{P}$ for the $x$ dimension, in the scenario where via-points are added after $t=1.0$. The different line colors correspond to different sigmoid function parameters $c_{1}$ and $c_{2}$.
A-G Analysis of the effect of $\delta_{\mathrm{al}}$, $\delta_{\mathrm{ep}}$.
Finally, Fig. 19 shows the stiffness profiles for changing values of $\delta_{\mathrm{al}}$, $\delta_{\mathrm{ep}}$, using our approach (5) in the scenario with via-points. In Fig. 19–left we see that increasing $\delta_{\mathrm{al}}$ has the effect of decreasing the overall stiffness in the regions where training data is present. This is due to $\delta_{\mathrm{al}}$ enlarging the predicted covariance $\bm{\Sigma}_{\mathrm{al}}^{*}$. Similarly, smaller values of $\delta_{\mathrm{ep}}$ (e.g., $\delta_{\mathrm{ep}}=1.0$, which directly applies the epistemic uncertainty from (4) to compute stiffness) may not sufficiently amplify the uncertainty to ensure that stiffness approaches zero outside the training data (e.g. the time instant just after $t=1.0$ in Fig. 7(b)–right). This reinforces our argument for introducing scaling parameters for the uncertainties, which are not tied to the KMP model’s data fit but are instead focused on shaping the desired behavior of the system in leveraging these uncertainties.
Figure 19: Stiffness values $\bm{G}_{P}$ for the $x$ dimension, in the scenario where via-points are added after $t=1.0$. The different line colors correspond to different epistemic and aleatoric scaling factors $\gamma_{\mathrm{al}}$ and $\gamma_{\mathrm{al}}$.

### v3

@@ I Introduction
We argue that in order to modulate skills locally, we require an underlying skill representation that 1) encodes trajectory distributions with both aleatoric and epistemic uncertainties—randomness in data and model knowledge gaps [7], and 2) allows for trajectory modulation trivially when modifications of the original demonstrations are required. To achieve this we build on the kernelized movement primitives (KMP) framework [8] (see Sec. III for a background review). Although task-parameterization of KMPs is briefly introduced in [8], the adaptation of local models has received little attention. We address the adaptation of local models by investigating when, where and how to add via-points. The result is an interactive learning framework of task-parameterized skills with local trajectory modulation (Sec. IV) that 1) permits the interactive correction of model imperfections locally, such that corrections ‘move’ with objects, 2) allows the definition of new task parameters interactively, when tasks change to require new/different objects and 3) ensures compliant interaction when extending skills beyond their initial duration by regulating stiffness based on the epistemic uncertainty, while simultaneously adding via-points. Figure 2 provides an overview of our approach. Our experimental evaluation (Sec. V) shows that our framework permits users to incrementally build on an initial model of a skill by interactively correcting errors and adding new behaviors in any task phase. We discuss the results and provide conclusions in Sec. VIVII. An overview of the notations and acronyms used is provided in the Additional Material B-AB-B.
@@ II-C Our approach and contribution
In Table I we compare our approach to other task-parameterized approaches on feature level. As shown in [5] task-parameterized approaches are more suitable for adapting to new situations than non-task-parameterized approaches like [30] or [31], including those involving multiple objects. To the best of our knowledge, we present the first task-parameterized approach allowing skill extensions with new frames and the application of local via-points, making it possible to move corrections with their corresponding objects and allowing interactive modulations, where corrections/extensions can be provided and incorporated into the policy online during execution.
@@ IV-D Uncertainty-aware skill extension in regions without demonstrations
Popular variable impedance schemes found in LfD regulate the robot stiffness by the inverse of covariance matrices [28, 23]. When the latter represent the aleatoric uncertainty, they provide an efficient way for robots to be more precise, by being stiffer, where demonstrations showed less variance, following a minimum intervention control principle [33, 28]. In the case of epistemic uncertainty, such schemes contribute to better compliance when the robot is uncertain about its actions [23]. In kernel-based methods, kernel hyperparameters are typically optimized for the training data, but their choice influences the behavior of the model in regions where data was not shown. For instance, the kernel length depends on the scale of the input domain, but it also dictates how quickly the epistemic uncertainty increases when moving away from the training data. [^2] Similar arguments can be made for other hyperparameters such as the noise variance, see [23]. Having ways to clearly distinguish between the two types of uncertainty enables users to better design variable impedance strategies. Similarly to [34], the covariance prediction of KMP can be decomposed into two terms, corresponding to aleatoric and epistemic components (see Appendix A-1 for the derivation):
@@
$\underbrace{\bm{k}^{**}-\bm{k}^{*}\bm{K}^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{ep}}^{*}}+\underbrace{\bm{k}^{*}\left(\bm{K}+\bm{K}(\lambda_{2}\bm{\Sigma})^{-1}\bm{K}\right)^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{al}}^{*}}$ (4)
@@ V-A Evaluation on real robot
We evaluate our approach on a torque-controlled 7-DoF robot in an industrial scenario where an inner ring of a ball bearing needs to be transferred between two boxes (‘box 1’ and ‘box 2’) placed at different locations on a workbench. [^4] You can find more information and videos at github.com/DLR-RM/interactive-incremental-learning We provide $M=4$ demonstrations with different box positions, see Fig. 1-top. We use a time-driven representation with $\bm{s}_{h,m}=t_{h,m}/T_{m}$, where $t$ is a time step, and learn the end-effector position $\bm{\xi}_{h,m}=\bm{x}_{h,m}$. In order to easily re-scale the skill duration, we map all the inputs to the interval $[0,1]$ by dividing them by the duration of each demonstration $T_{m}$. The experiments start with $P=2$, with task parameters $\bm{b}^{(p)},\bm{A}^{(p)}$ representing the box positions and orientations, respectively. All local KMPs were initialized from GMMs with 12 components, trained on locally projected data, and $N=500$ inputs, equally spread over the input space. We chose a Matérn kernel ($\nu=5/2$) with length scale $l=0.1$ and noise variance 1.0 (see [35]). Other KMP hyperparameters were $\lambda_{1}=0.1$, $\lambda_{2}=1$, $\alpha=1$, chosen empirically. For completeness, we provide an overview of hyperparameters in Appendix A-2. In all experiments, via-points are added with $\gamma_{\Sigma}=1\times 10^{-8}$. Our algorithm successfully generalizes to novel box positions. However, when they deviate significantly from the demonstrated ones, TP-GMM [2], as well as the original TP-KMP formulation [8] fail (see Fig. 1-bottom).
@@ V-B Experiment 1: Improving generalization with interactively-defined local via-points
Since all demonstrations are given at the same height, the model has a high confidence for the $z$ coordinate in all frames, which leads to poor generalization when one of the boxes is moved to a different height (Fig. 1-bottom). If box 2 is moved to a lower height, the trade-off found by the model, from the expected vertical motions in the two frames, is not high enough to successfully move the ring out of box 1, leading to a collision. Using Algorithm 1, we have the ability to directly correct the robot and set via-points based on corrections made online. In this experiment we used an external force trigger, with threshold $\gamma_{F}=20N$. Figure 3 shows the results obtained for this experiment. With the robot running the model, a user applies a force at the robot end-effector, helping it avoid the collision and successfully move the ring out of the box. Via-points are added to the frame of box 1 and, through $\bm{b}^{(1)},\bm{A}^{(1)}$, are mapped to the global frame. With the added via-points, the robot achieves a success rate of $93\%$ across 15 box configurations (with three different heights), improved over the $60\%$ success rate of TP-GMM [2] and original TP-KMP [8] baselines (see Additional Material B-C).
@@ V-D Experiment 3: Incremental learning at non-demonstrated inputs
We consider a baseline where the gains are computed from the original KMP covariance (3) as $\bm{G}_{P}=(\mathrm{cov}\left[\bm{\xi}(\bm{s}^{*})\right])^{-1}$. To address numerical issues and ensure a maximum stiffness value we regularize $\mathrm{cov}\left[\bm{\xi}(\bm{s}^{*})\right]$, $\delta_{\mathrm{al}}\bm{\Sigma}^{*}_{\mathrm{al}}$, $\delta_{\mathrm{ep}}\bm{\Sigma}^{*}_{\mathrm{ep}}$ with a small scalar $1.5\times 10^{-3}$. Figure 5 shows that using our approach (5) the interaction forces after ${t=1.0}$ are negligible. Note that far from the training data, the KMP expectation (2) converges to zero, as we assume a zero mean prior (similarly to GPs [35]). With the improved compliance behaviors introduced by (5), the user is able to add via-points such that the robot moves up after the ring insertion. In this experiment we used a distance threshold with $\gamma_{\xi}=0.2m$, taking advantage of the reference trajectory going to zero to trigger the definition of via-points. The bottom-left image in Fig. 6 shows the increase in distance as via-points are added, while the plots on the right show the resulting robot motion and uncertainties. Appendix A-3 shows the detailed stiffness profiles from Experiment 3 including a sensitivity analysis of the kernel length scales.
@@ Acknowledgments
The authors thank Thomas Eiband and Korbinian Nottensteiner for their discussions.
@@ Appendix A Additional material
A-1 Decomposition of KMP covariance into a sum of epistemic and aleatoric terms
$\mathrm{cov}\left[\bm{\xi}(\bm{s}^{*})\right]=$
$=\bm{k}^{**}-\bm{k}^{*}\left[\bm{K}^{-1}-\bm{K}^{-1}\left(\bm{K}^{-1}+(\lambda_{2}\bm{\Sigma})^{-1}\right)^{-1}\bm{K}^{-1}\right]\bm{k}^{*\top}$
$=\bm{k}^{**}-\bm{k}^{*}\bm{K}^{-1}\bm{k}^{*\top}+\bm{k}^{*}\bm{K}^{-1}\left(\bm{K}^{-1}+(\lambda_{2}\bm{\Sigma})^{-1}\right)^{-1}\bm{K}^{-1}\bm{k}^{*\top}$
$=\underbrace{\bm{k}^{**}-\bm{k}^{*}\bm{K}^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{ep}}^{*}}+\underbrace{\bm{k}^{*}\left(\bm{K}+\bm{K}(\lambda_{2}\bm{\Sigma})^{-1}\bm{K}\right)^{-1}\bm{k}^{*\top}}_{\bm{\Sigma}_{\mathrm{al}}^{*}},$
A-2 Hyper-parameter overview
TABLE III: Hyper-parameter overview
Hyper-parameter Used in paper Validated range/options Remark
Scaling factor for KMP covariance $\alpha$ 1.0 [0.1; 1.0] -
Kernel function $k(.,.)$ matern2 matern2 / rbf -
between aleatoric and epistemic uncertainty
Table III shows the different hyperparameters used in this paper and a tested range of values to provide a guideline for choosing hyperparameters.
A-3 Detailed stiffness profiles for Experiment 3 with kernel length scale sensitivity analysis
Appendix B Additional material
B-A Key notations
Table IV summarizes the key notations used in our framework.
TABLE IV: Description of key notations
@@ B-A Key notations
B-B Key acronyms
Table V shows a glossary of important acronyms in our approach.
TABLE V: Glossary of important acronyms
@@ B-B Key acronyms
B-C Evaluation details
Table VI and Table VII show more details of the quantitative evaluation of 15 different task-parameterized scenarios, mentioned in Sec. V-B of the paper.